# Fission mode identification in the 180Hg region: derivative analysis approach

### 17 Mar 2026 v1

@@ Introduction
The process of nuclear fission, i.e. the splitting of a single nucleus into several fragments, remains one of the least comprehended fundamental phenomena in nuclear physics. Although the macroscopic Liquid Drop Model (LDM) could quantitatively explain the phenomenon of fission straight after its discovery in 1938 [14], its full understanding is yet to be achieved. The historical observation of unequal mass split in 236U, known today as the asymmetric-fission mode ($A$ mode), has been confirmed in various experiments with different actinide nuclei [14]. This $A$ mode produces a double-humped fission-fragment mass distribution (FFMD) and is understood to be driven by the nucleon pairing and shell effects. The latest theoretical works refer to the octupole deformation of fragments with atomic numbers $Z=52$ and 56 [20, 21] as responsible for the major mass-asymmetric splits in fission of actinide nuclei. In contrast, the symmetric-fission mode ($S$ mode) produces fragments with comparable masses and thus the relevant FFMD exhibits a Gaussian-like shape. The other global difference between the $A$ and $S$ modes is the characteristic values of the total kinetic energy (TKE) of fission products known to be lower for the $S$ mode than for the $A$-mode. This difference is understood as the consequence of more elongated fission configuration at scission point.
@@
Here, the functions $f^{{}^{\prime}}(A)$ and $f^{{}^{\prime\prime}}(A)$ are composed of $i$ components relevant to each peak in the corresponding fission mode, with $B_{i}$, $\sigma_{A_{i}}$ and $\overline{A_{i}}$ being the mode amplitude, width and peak position in the FFMD, respectively. The number of Gaussians to use is given by the number $i$ of minima identified in the 2nd derivative data. For example, ii=5 would signify five components in the FFMD, one for the $S$-mode, two for the first $A$-mode ($A1$) and further two for the second $A$-mode ($A2$). The term $f_{i}(A)$ describing the $S$- and $A$-mode components in the FFMD is present in all the three expressions of Eq.1. This means that, once the composition of the fit function is determined, the FFMD composition can be obtained from the fit made on any data set (FFMD, 1st or 2nd derivative).
However, it appears that the use of Eq.1 with completely uncorrelated and unconstrained parameters may provide good overall data description in terms of $\chi^{2}$ but gives meaningless results for fission modes: the $S$-mode position may appear anywhere in the FFMD or non-identical integral area obtained for the same $A$-mode components. Therefore, it is recommended to fix the $S$-mode position at $\frac{{A_{CN}}}{2}$ and to demand identical widths, amplitudes as well as displacement from the $S$-mode position for the same $A$-mode components.
@@ Results
Simulated data: 240Pu. The FFMD of a nucleus in the classical actinide region can be best described by using one $S$- and three $A$-fission modes; here, they are denoted using the Brosa nomenclature [3] as $SL$ and SS1-SS3, correspondingly. Each mode is characterised by its position in the FFMD relevant to its origin, by its specific characteristic average TKE ($\overline{TKE}$) relevant to the scission configuration, and by its relative weight $W$ indicating the mode importance in the considered FFMD.
@@
Presence of 7 minima is clearly seen in the simulated data set. This number determines composition of the fit function and implies presence of one $S$-mode and three $A$-modes in the considered FFMD, in accordance with the input data.
@@
From the one- (1D) and two-dimensional (2D) fits with the 7G fit function on the simulated and resolution affected data, the $S$- and $A$-mode parameters are extracted and parameters of the $S$- and $A$-modes are obtained and can subsequently be compared with the input values. Table 1 demonstrates this sort of comparison made for a range of $E^{*}$ from 10 to 70 MeV. The following can be stated:
  • For $E^{*}\geq$30 MeV, a direct FFMD fit (1D in Table 1) does not correctly reproduce either the mode positions or their respective weights, the disagreement with input data being sometimes stronger for the resolution-affected data. The striking feature here is the suppression of the $SL$-mode, in favour of the $S1$-mode.
@@
Simulated data: A=180. The FFMD of this hypothetical nucleus can be simulated by assuming different $S$ and $A$ mode combinations. In the simulation, we focus on the two selected combinations: one $S$ plus one $A$ mode and one $S$ plus two $A$ modes, the $A$ modes displaced by 10 and 24 u with respect to the centroid $\frac{A_{CN}}{2}=90$.
@@
As seen above, the resolution effect manifests as a flattening of structural features in the FFMD and its derivative data. The right column of Fig.2 demonstrates the power of the $2^{nd}$ derivative analysis: it appears possible to identify presence of $A$ modes in the FFMDs strongly dominated by the $S$ mode, cf. panels (f,j). In these extreme cases, the presence of the $A1$ mode is evidenced either by a shallow minimum at the $\frac{A_{CN}}{2}$, if compared to the pure $S$-mode derivative from panel (l). The $A1$ mode position can in this case be fixed and fitted from the minima of the more structured FFMDs, such as in panels (b,d), where the position can be read directly by eye. Panel (j) also illustrates the sensibility of the method to detect the $A2$ contribution to the FFMD, which is kept here on the detection limit. The minimum visible at $A\sim$66 is expected to become flatter or be completely blurred if the data are measured with resolution worse than 2 u and when the $A2$ mode contribution is less than 4%, similar to the $S3$ mode in Fig.1(b).
To clarify on how the choice of the fitting function may affect the determination of the major fission-mode parameters, the 3G and 5G functions were used to fit data from Fig.2. For each fit, the $S$ mode position was kept fixed at 90 u, whereas the $A$ modes parameters were kept free but identical for each particular $A$ mode. Table 2 compares the obtained results with the input data.
@@
The FFMDs have consequently been fitted with the 3G function. The fits were made with both free and fixed position of the $A$ mode, cf. Table 3 and Fig.3 for the results. A juxtaposition of these results gives rise to the following observations:
  • FFMD fit with free $A$-mode parameters shows a drift of the mode position with $E^{*}$, in contrast to the 2nd derivative data, as well as an irregular change in width $\sigma$ and weight $W$. Conversely, the FFMD fit with fixed $A$-mode positions results in increasing $A$- and $S$-mode weights $W$ and widths $\sigma$ as a function of $E^{*}$, a trend that seems physically plausible. Moreover, such a fit does not lead to the loss of accuracy in the data description, cf. reduced $\chi^{2}$ values in Table 3.
@@
Figure 3: Experimental FFMD for 180H at two different excitation energies taken from Ref. [17] and their derivatives calculated with window length of 3 u (panels (b,f)) and 5 u (panels (d,h)). The minima identified at AA=78, 90 and 102 u in the 2nd derivative data determine the fit function composition as $S$ plus one AA mode. Panels (a,e) and (c,g) display, respectively, the individual mode components from the FFMDs fit with free and fixed $A$-mode peak positions. Blue/green lines indicate the SS/AA-mode components, red line is their sum. The mass resolution was 2.4 u.
@@ Conclusions
Furthermore, the direct FFMD fits of the AA=180 nucleus using 3G and 5G fitting functions show that, without a priori knowledge of the appropriate functional composition, the extracted results may lead to incorrect physical results: increase of the $A$-mode contribution with increasing excitation energy.
@@
If applied to real experimental data (180Hg from Ref. [17]), the 2nd derivative analysis shows itself capable of detecting, in a simple and undoubted way, the presence of both $S$ and $A$ modes in the FFMD data, thus resolving the question on the $S$-mode presence in the data raised by the authors. The $A$-mode weight extracted with a fixed $A$-mode position exhibits the anticipated increase with excitation energy, lending additional credibility to the analysis method applied to datasets with limited statistics.

### 19 Mar 2026 v2

@@ Introduction
The process of nuclear fission, i.e. the splitting of a single nucleus into several fragments, remains one of the least comprehended fundamental phenomena in nuclear physics. Although the macroscopic Liquid Drop Model (LDM) could quantitatively explain the phenomenon of fission straight after its discovery in 1938 [14], its full understanding is yet to be achieved. The historical observation of unequal mass split in 236U, known today as the asymmetric-fission mode (AA mode), has been confirmed in various experiments with different actinide nuclei [14]. This AA mode produces a double-humped fission-fragment mass distribution (FFMD) and is understood to be driven by the nucleon pairing and shell effects. The latest theoretical works refer to the octupole deformation of fragments with atomic numbers $Z=52$ and 56 [20, 21] as responsible for the major mass-asymmetric splits in fission of actinide nuclei. In contrast, the symmetric-fission mode ($S$ mode) produces fragments with comparable masses and thus the relevant FFMD exhibits a Gaussian-like shape. The other global difference between the AA and $S$ modes is the characteristic values of the total kinetic energy (TKE) of fission products known to be lower for the $S$ mode than for the AA-mode. This difference is understood as the consequence of more elongated fission configuration at scission point.
@@
Here, the functions $f^{{}^{\prime}}(A)$ and $f^{{}^{\prime\prime}}(A)$ are composed of $i$ components relevant to each peak in the corresponding fission mode, with $B_{i}$, $\sigma_{A_{i}}$ and $\overline{A_{i}}$ being the mode amplitude, width and peak position in the FFMD, respectively. The number of Gaussians to use is given by the number $i$ of minima identified in the 2nd derivative data. For example, ii=5 would signify five components in the FFMD, one for the $S$-mode, two for the first AA-mode ($A1$) and further two for the second AA-mode ($A2$). The term $f_{i}(A)$ describing the $S$- and AA-mode components in the FFMD is present in all the three expressions of Eq.1. This means that, once the composition of the fit function is determined, the FFMD composition can be obtained from the fit made on any data set (FFMD, 1st or 2nd derivative).
However, it appears that the use of Eq.1 with completely uncorrelated and unconstrained parameters may provide good overall data description in terms of $\chi^{2}$ but gives meaningless results for fission modes: the $S$-mode position may appear anywhere in the FFMD or non-identical integral area obtained for the same AA-mode components. Therefore, it is recommended to fix the $S$-mode position at $\frac{{A_{CN}}}{2}$ and to demand identical widths, amplitudes as well as displacement from the $S$-mode position for the same AA-mode components.
@@ Results
Simulated data: 240Pu. The FFMD of a nucleus in the classical actinide region can be best described by using one $S$- and three AA-fission modes; here, they are denoted using the Brosa nomenclature [3] as SLSL and SS1-SS3, correspondingly. Each mode is characterised by its position in the FFMD relevant to its origin, by its specific characteristic average TKE ($\overline{TKE}$) relevant to the scission configuration, and by its relative weight $W$ indicating the mode importance in the considered FFMD.
@@
Presence of 7 minima is clearly seen in the simulated data set. This number determines composition of the fit function and implies presence of one $S$-mode and three AA-modes in the considered FFMD, in accordance with the input data.
@@
From the one- (1D) and two-dimensional (2D) fits with the 7G fit function on the simulated and resolution affected data, the $S$- and AA-mode parameters are extracted and parameters of the $S$- and AA-modes are obtained and can subsequently be compared with the input values. Table 1 demonstrates this sort of comparison made for a range of $E^{*}$ from 10 to 70 MeV. The following can be stated:
  • For $E^{*}\geq$30 MeV, a direct FFMD fit (1D in Table 1) does not correctly reproduce either the mode positions or their respective weights, the disagreement with input data being sometimes stronger for the resolution-affected data. The striking feature here is the suppression of the SLSL-mode, in favour of the $S1$-mode.
@@
Simulated data: A=180. The FFMD of this hypothetical nucleus can be simulated by assuming different $S$ and AA mode combinations. In the simulation, we focus on the two selected combinations: one $S$ plus one AA mode and one $S$ plus two AA modes, the AA modes displaced by 10 and 24 u with respect to the centroid $\frac{A_{CN}}{2}=90$.
@@
As seen above, the resolution effect manifests as a flattening of structural features in the FFMD and its derivative data. The right column of Fig.2 demonstrates the power of the $2^{nd}$ derivative analysis: it appears possible to identify presence of AA modes in the FFMDs strongly dominated by the $S$ mode, cf. panels (f,j). In these extreme cases, the presence of the $A1$ mode is evidenced either by a shallow minimum at the $\frac{A_{CN}}{2}$, if compared to the pure $S$-mode derivative from panel (l). The $A1$ mode position can in this case be fixed and fitted from the minima of the more structured FFMDs, such as in panels (b,d), where the position can be read directly by eye. Panel (j) also illustrates the sensibility of the method to detect the $A2$ contribution to the FFMD, which is kept here on the detection limit. The minimum visible at $A\sim$66 is expected to become flatter or be completely blurred if the data are measured with resolution worse than 2 u and when the $A2$ mode contribution is less than 4%, similar to the $S3$ mode in Fig.1(b).
To clarify on how the choice of the fitting function may affect the determination of the major fission-mode parameters, the 3G and 5G functions were used to fit data from Fig.2. For each fit, the $S$ mode position was kept fixed at 90 u, whereas the AA modes parameters were kept free but identical for each particular AA mode. Table 2 compares the obtained results with the input data.
@@
The FFMDs have consequently been fitted with the 3G function. The fits were made with both free and fixed position of the AA mode, cf. Table 3 and Fig.3 for the results. A juxtaposition of these results gives rise to the following observations:
  • FFMD fit with free AA-mode parameters shows a drift of the mode position with $E^{*}$, in contrast to the 2nd derivative data, as well as an irregular change in width $\sigma$ and weight $W$. Conversely, the FFMD fit with fixed AA-mode positions results in increasing AA- and $S$-mode weights $W$ and widths $\sigma$ as a function of $E^{*}$, a trend that seems physically plausible. Moreover, such a fit does not lead to the loss of accuracy in the data description, cf. reduced $\chi^{2}$ values in Table 3.
@@
Figure 3: Experimental FFMD for 180H at two different excitation energies taken from Ref. [17] and their derivatives calculated with window length of 3 u (panels (b,f)) and 5 u (panels (d,h)). The minima identified at AA=78, 90 and 102 u in the 2nd derivative data determine the fit function composition as $S$ plus one AA mode. Panels (a,e) and (c,g) display, respectively, the individual mode components from the FFMDs fit with free and fixed AA-mode peak positions. Blue/green lines indicate the SS/AA-mode components, red line is their sum. The mass resolution was 2.4 u.
@@ Conclusions
Furthermore, the direct FFMD fits of the AA=180 nucleus using 3G and 5G fitting functions show that, without a priori knowledge of the appropriate functional composition, the extracted results may lead to incorrect physical results: increase of the AA-mode contribution with increasing excitation energy.
@@
If applied to real experimental data (180Hg from Ref. [17]), the 2nd derivative analysis shows itself capable of detecting, in a simple and undoubted way, the presence of both $S$ and AA modes in the FFMD data, thus resolving the question on the $S$-mode presence in the data raised by the authors. The AA-mode weight extracted with a fixed AA-mode position exhibits the anticipated increase with excitation energy, lending additional credibility to the analysis method applied to datasets with limited statistics.